# Analytic action principle in ${\cal N}=2$ AdS4 harmonic superspace

### v1

@@ 4 Analytic action principle
$\int d^{4|8}z\,E\,{\cal D}^{i}_{\alpha}\Psi_{i}^{\alpha}=\int d^{4|8}z\,E\,\bar{\cal D}_{i}^{\dot{\alpha}}\bar{\Psi}^{i}_{\dot{\alpha}}=0\,$, (4.2)
the action (4.1) may be expressed via the analytic projector (3.10):
$S=$ $\frac{1}{16}\int d^{4|8}z\,E\int du\frac{1}{(S^{++})^{2}}[({\cal D}^{+})^{2}+4S^{++}][(\bar{\cal D}^{+})^{2}+4S^{++}]{\cal L}$
$=$ $\int d^{4|8}z\,E\int du\frac{{\cal L}^{(+4)}}{(S^{++})^{2}}\,$, (4.3)
@@
$S=\int d^{4|8}z\,E\int du\frac{U}{\Pi^{(+4)}U}{\cal L}^{(+4)}\,$, (4.4)
where $U$ is an arbitrary superfield such that $\Pi^{(+4)}U$ is nowhere vanishing. Integrating by parts covariant derivatives, it is possible to show that this action is independent of a particular choice of $U$, $\delta_{U}S=0$. The action (4.4) reduces to (4.3) for $U=1$.
[…]
$S=\frac{1}{2\pi}\oint(u^{+}du^{+})\int d^{4|8}z\,E\frac{{\cal L}^{++}}{(S^{++})^{2}}\,.$ (4.5)
Unlike the projective superspace approach, in the harmonic superspace one, the factor $(S^{++})^{-2}$ in the action (4.3) may represent a potential problem because of harmonic singularities. It is a non-trivial exercise to check that this factor cancels out upon calculation of the component structure of this action. As we will show below, there is an alternative action principle for an analytic Lagrangian where this problem does not occur.
Consider an analytic superfield Lagrangian ${\cal L}^{(+4)}$, ${\cal D}^{+}_{\alpha}{\cal L}^{(+4)}=\bar{\cal D}^{+}_{\dot{\alpha}}{\cal L}^{(+4)}=0$. By analogy with the AdS5 harmonic superspace considered in Refs. [22, 23], we will look for the invariant action for ${\cal L}^{(+4)}$ in the following form:
$S=$ $S_{0}+aS_{1}\,$, (4.6a)
$S_{0}=$ $\int d^{4}x\,e\int du({\cal D}^{-})^{4}{\cal L}^{(+4)}|\,$, (4.6b)
$S_{1}=$ $\int d^{4}x\,e\int du\,S^{--}[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}]{\cal L}^{(+4)}\big|\,$, (4.6c)
@@
$({\cal D}^{-})^{4}:=\frac{1}{16}({\cal D}^{-})^{2}(\bar{\cal D}^{-})^{2}\,.$ (4.7)
The coefficient $a$ in (4.6a) is to be found from the condition of invariance of this action under the isometry of AdS4|8, $\delta_{\xi}S=0$. Since the Lagrangian is a scalar, it transforms under (2.15) as
$\delta_{\xi}{\cal L}^{(+4)}=(\xi+\rho{\cal J}){\cal L}^{(+4)}\,.$ (4.8)
@@
$\int d^{4}x\,e\int du\,{\cal D}_{a}\Phi^{a}\big|=$ $0\,$, (4.9a)
$\int d^{4}x\,e\int du\,D^{++}\Phi^{--}\big|=\int d^{4}x\,e\int du\,D^{--}\Phi^{++}\big|=$ $0\,$, (4.9b)
$\int d^{4}x\,e\int du\,{\cal J}\Phi^{(0)}\big|=$ $0\,$, (4.9c)
we can represent the variations of $S_{0}$ and $S_{1}$ in the forms
$\delta_{\xi}S_{0}=$ $\int d^{4}x\,e\int du\left(-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\right)({\cal D}^{-})^{4}{\cal L}^{(+4)}\big|$ (4.10a)
$\delta_{\xi}S_{1}=$ $\int d^{4}x\,e\int du\,S^{--}\left(-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\right)[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}]{\cal L}^{(+4)}\big|\,.$ (4.10b)
The next procedure consists in the following steps:
@@
    Commute the derivatives ${\cal D}^{+}_{\alpha}$ and $\bar{\cal D}^{+\dot{\alpha}}$ with $({\cal D}^{-})^{2}$ and $(\bar{\cal D}^{-})^{2}$ using the algebra of covariant derivatives (3.7) and the analyticity of the Lagrangian, ${\cal D}^{+}_{\alpha}{\cal L}^{(+4)}=\bar{\cal D}^{+\dot{\alpha}}{\cal L}^{(+4)}=0$.
@@
    Using the same algebra of covariant derivatives, in the integrand pull the operators ${\cal D}_{a}$ and $\cal J$ to the left and use the rules of integration by parts (4.9).
@@
    $\sigma^{a}_{\alpha\dot{\alpha}}{\cal D}_{a}\xi^{-\alpha}=$ $2i(S^{--}\bar{\xi}^{+}_{\dot{\alpha}}-S^{+-}\bar{\xi}^{-}_{\dot{\alpha}})\,$, (4.11)
    $(\sigma^{a})^{\alpha\dot{\alpha}}{\cal D}_{a}\bar{\xi}^{-}_{\dot{\alpha}}=$ $-2i(S^{--}\xi^{+\alpha}-S^{+-}\xi^{-\alpha})\,.$ (4.12)
    These identities follow from Eq. (2.11c).
As a result, we find the following expressions for the variations (4.10):
$\delta_{\xi}S_{0}=$ $\frac{1}{2}\int d^{4}x\,e\int du\,S^{--}\Big[\xi^{+\alpha}{\cal D}^{-}_{\alpha}(\bar{\cal D}^{-})^{2}-\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}({\cal D}^{-})^{2}-8S^{--}\xi^{+\alpha}{\cal D}^{-}_{\alpha}+8S^{--}\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}$
$+12S^{+-}\xi^{-\alpha}{\cal D}^{-}_{\alpha}-12S^{+-}\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\Big]{\cal L}^{(+4)}\big|\,$, (4.13)
$\delta_{\xi}S_{1}=$ $-\int d^{4}x\,e\int du\,S^{--}\Big[\xi^{+\alpha}{\cal D}^{-}_{\alpha}(\bar{\cal D}^{-})^{2}-\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}({\cal D}^{-})^{2}-8S^{--}\xi^{+\alpha}{\cal D}^{-}_{\alpha}+8S^{--}\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}$
$+12S^{+-}\xi^{-\alpha}{\cal D}^{-}_{\alpha}-12S^{+-}\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\Big]{\cal L}^{(+4)}\big|\,.$ (4.14)
Hence, for $a=1/2$, the action (4.6a) is invariant under the isometry of AdS4|8, $\delta_{\xi}(S_{0}+S_{1})=0$. Finally, we write the action principle for an analytic Lagrangian ${\cal L}^{(+4)}$ in the form:
$\framebox{$\displaystyle S=\int d^{4}x\,e\int du\left\{({\cal D}^{-})^{4}+\frac{1}{2}S^{--}\left[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}\right]\right\}{\cal L}^{(+4)}\bigg|\,.$}$ (4.15)
In this form, the analytic action principle is free from any harmonic singularities and is suitable for studying the component structure of the analytic Lagrangian ${\cal L}^{(+4)}$. Below, we demonstrate how the analytic action principle (4.15) applies for calculation of the component structure in the free $q$-hypermultiplet model.
5 Example: component structure of the free hypermultiplet action
@@
${\cal L}^{(+4)}=-\tilde{q}^{+}D^{++}q^{+}\,$, (5.1)
where $q^{+}$ is an an analytic superfield and $\tilde{q}^{+}$ is its tilde-conjugate,
${\cal D}^{+}_{\alpha}q^{+}=\bar{\cal D}^{+}_{\dot{\alpha}}q^{+}=0\,.$ (5.2)
Note that this superfield is not required to be annihilated by $D^{++}$.
To evaluate the component structure of the action (4.15) with the Lagrangian (5.1), we proceed by the following steps:
@@
    We commute the derivatives $\mathcal{D}^{-}_{\alpha}$ and $\bar{\mathcal{D}}^{-\dot{\alpha}}$ to the right and act on $q^{+}$ and $\tilde{q}^{+}$. The commutators ${\cal D}^{-}_{\alpha},D^{++}]=-{\cal D}^{+}_{\alpha}$ and $\bar{\cal D}^{-\dot{\alpha}},D^{++}]=-\bar{\cal D}^{+\dot{\alpha}}$ produce the derivatives $\mathcal{D}^{+}$’s, which annihilate the analytic superfield $q^{+}$ according to Eq. (5.2).
@@
    The operators ${\cal D}_{a}$ and $\cal J$ are moved to the left by using the rules for integration by parts (4.9).
@@
$S_{q}=$ $S_{\mathrm{bos}}+S_{\mathrm{ferm}}\,$, (5.3)
$S_{\mathrm{bos}}=$ $-\int d^{4}x\,e\int du$
$\times\Bigl[\bigl((\mathcal{D}^{-})^{4}\tilde{q}^{+}\bigr)D^{++}q^{+}+\tilde{q}^{+}D^{++}\bigl((\mathcal{D}^{-})^{4}q^{+}\bigr)+\frac{1}{4}(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}\tilde{q}^{+})D^{++}(\mathcal{D}^{-}_{\beta}\bar{\mathcal{D}}^{-}_{\dot{\alpha}}q^{+})$
@@
$+\frac{i}{2}(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}\tilde{q}^{+})(\mathcal{D}_{\beta\dot{\alpha}}q^{+})-\frac{i}{2}(\mathcal{D}_{\beta\dot{\alpha}}\tilde{q}^{+})(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}q^{+})$
$+\frac{1}{4}\bigl((\mathcal{D}^{-})^{2}\tilde{q}^{+}+(\bar{\mathcal{D}}^{-})^{2}\tilde{q}^{+}\bigr)(\mathcal{J}+2S^{--}D^{++})q^{+}$
$+\frac{1}{4}\tilde{q}^{+}(\mathcal{J}+2D^{++}S^{--})\bigl((\mathcal{D}^{-})^{2}q^{+}+(\bar{\mathcal{D}}^{-})^{2}q^{+}\bigr)+9S^{--}\tilde{q}^{+}\mathcal{J}q^{+}\Bigr]\bigg|\,$, (5.4)
$S_{\mathrm{ferm}}=$ $-\int d^{4}x\,e\int du$
$\times\Bigl[\frac{1}{8}(\mathcal{D}^{-\alpha}(\bar{\mathcal{D}}^{-})^{2}\tilde{q}^{+})D^{++}(\mathcal{D}^{-}_{\alpha}q^{+})+\frac{1}{8}(\bar{\mathcal{D}}^{-}_{\dot{\beta}}\tilde{q}^{+})D^{++}(\bar{\mathcal{D}}^{-\dot{\beta}}(\mathcal{D}^{-})^{2}q^{+})$
@@
$+\frac{i}{2}(\mathcal{D}^{-\alpha}\tilde{q}^{+})(\mathcal{D}_{\alpha\dot{\beta}}\bar{\mathcal{D}}^{-\dot{\beta}}q^{+})-\frac{i}{2}(\bar{\mathcal{D}}^{-\dot{\beta}}\tilde{q}^{+})(\mathcal{D}_{\alpha\dot{\beta}}\mathcal{D}^{-\alpha}q^{+})$
$+\frac{1}{2}(\mathcal{D}^{-\alpha}\tilde{q}^{+})(\mathcal{J}+2S^{+-}+2S^{--}D^{++})(\mathcal{D}^{-}_{\alpha}q^{+})$
$+\frac{1}{2}(\bar{\mathcal{D}}^{-}_{\dot{\beta}}\tilde{q}^{+})(\mathcal{J}+2S^{+-}+2S^{--}D^{++})(\bar{\mathcal{D}}^{-\dot{\beta}}q^{+})\Bigr]\bigg|\,.$ (5.5)
@@
${\cal D}^{-}_{\beta}\bar{\cal D}^{-}_{\dot{\alpha}}q^{+}|=$ $-i\sigma^{a}_{\beta\dot{\alpha}}A^{-}_{a}(x,u)=-iA_{\beta\dot{\alpha}}^{-}(x,u)\,$, (5.6f)
@@
These components enter the actions (5.4) and (5.5) as follows:
@@
$-\frac{1}{2}\Phi^{\alpha}({\cal J}+2S^{+-}+2S^{--}D^{++})\Psi_{\alpha}+\frac{1}{2}\bar{\Psi}_{\dot{\alpha}}({\cal J}+2S^{+-}+2S^{--}D^{++})\bar{\Phi}^{\dot{\alpha}}\Big].$ (5.7b)

### v2

@@ IV Analytic action principle
$\int d^{4|8}z\,E\,{\cal D}^{i}_{\alpha}\Psi_{i}^{\alpha}=\int d^{4|8}z\,E\,\bar{\cal D}_{i}^{\dot{\alpha}}\bar{\Psi}^{i}_{\dot{\alpha}}=0\,$, (35)
the action (34) may be expressed via the analytic projector (31):
$S=\int d^{4|8}z\,E\int du\frac{{\cal L}^{(+4)}}{(S^{++})^{2}}\,$, (36)
where
${\cal L}^{(+4)}=$ $\frac{1}{16}[({\cal D}^{+})^{2}+4S^{++}][(\bar{\cal D}^{+})^{2}+4S^{++}]{\cal L}$
$=$ $\Pi^{(+4)}{\cal L}$ (37)
is an analytic Lagrangian. It obeys the reality condition under the tilde-conjugation in the harmonic superspace [2], $\tilde{\cal L}^{(+4)}={\cal L}^{(+4)}$.
The analytic action principle in the form (36) was derived in the general setup of ${\cal N}=2$ supergravity matter theory in the harmonic superspace in Ref. [20]. It was proved there that this action is free from harmonic singularities for arbitrary covariantly analytic Lagrangian ${\cal L}^{(+4)}$ despite the fact that it contains the charged harmonic superfield $(S^{++})^{2}$ in the denominator. Here we apply the same action principle for matter theories on the particular AdS4|8 background with the covariantly constant torsion superfield $S^{ij}$.
$S=\int d^{4|8}z\,E\int du\frac{U}{\Pi^{(+4)}U}{\cal L}^{(+4)}\,$, (38)
where $U$ is an arbitrary nowhere vanishing superfield and $\Pi^{(+4)}U$ is its analytic projection. Integrating the covariant derivatives by parts, one can show that this action is independent of the particular choice of $U$,
$\delta_{U}S=0\,.$ (39)
Thus the superfield $U$ is pure gauge and the action (38) reduces to (36) for $U=1$. In Ref. [20], it was shown that this superfield $U$ originates from the vector multiplet compensator in the ${\cal N}=2$ supergravity. See also Ref. [34] for a review of covariant superspace approaches to ${\cal N}=2$ supergravity.
@@
$S=\frac{1}{2\pi}\oint(u^{+}du^{+})\int d^{4|8}z\,E\frac{{\cal L}^{++}}{(S^{++})^{2}}\,.$ (40)
Unlike the projective superspace approach, in the harmonic superspace one, the factor $(S^{++})^{-2}$ in the action (36) may represent a potential problem because of harmonic singularities. It is a non-trivial exercise to check that this factor cancels out upon calculation of the component structure of this action. As we will show below, there is an alternative action principle for an analytic Lagrangian where this problem does not occur.
Rather than further unfolding the action functional (36) we switch to a bottom-up approach: starting from the standard action principle in the flat harmonic superspace for an analytic Lagrangian [2], we seek for its direct AdS4 harmonic superspace generalization. Indeed, given an analytic Lagrangian $L^{(+4)}$ in the flat harmonic superspace, the corresponding action may be written as $S=\int d^{4}x\int du(D^{-})^{4}L^{(+4)}|$, with flat superspace derivatives $D^{-}_{\alpha},\bar{D}^{-}_{\dot{\alpha}}$ and ‘$|$’ standing for the $\theta=0$ projection. In the curved superspace, the flat superspace derivatives should be replaced with the covariant ones, ${\cal D}^{-}_{\alpha}$ and $\bar{\cal D}^{-}_{\dot{\alpha}}$, and the Lagrangian is represented by a covariantly analytic superfield ${\cal L}^{(+4)}$. Then, considering the mass dimension of torsion superfield $S^{--}$ and balance of the $U(1)$ charge, it is natural to search for an AdS4 generalization of the flat superspace action within the following most general ansatz:
$S=$ $S_{0}+a_{1}S_{1}+a_{2}S_{2}\,$, (41a)
$S_{0}=$ $\int d^{4}x\,e\int du({\cal D}^{-})^{4}{\cal L}^{(+4)}|\,$, (41b)
$S_{1}=$ $\int d^{4}x\,e\int du\,S^{--}[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}]{\cal L}^{(+4)}\big|\,$, (41c)
$S_{2}=$ $\int d^{4}x\,e\int du\,(S^{--})^{2}{\cal L}^{(+4)}|,$ (41d)
@@
$({\cal D}^{-})^{4}:=\frac{1}{16}({\cal D}^{-})^{2}(\bar{\cal D}^{-})^{2}\,.$ (42)
The coefficients $a_{1}$ and $a_{2}$ in (41a) are to be found from the condition of invariance of this action under the isometry of AdS4|8, $\delta_{\xi}S=0$.
Since the Lagrangian is a scalar, it transforms under (15) by the rule
$\delta_{\xi}{\cal L}^{(+4)}=(\xi+\rho{\cal J}){\cal L}^{(+4)}\,.$ (43)
@@
$\int d^{4}x\,e\int du\,{\cal D}_{a}\Phi^{a}\big|=$ $0\,$, (44a)
$\int d^{4}x\,e\int du\,D^{++}\Phi^{--}\big|=$ $0\,$, (44b)
$\int d^{4}x\,e\int du\,D^{--}\Phi^{++}\big|=$ $0\,$, (44c)
$\int d^{4}x\,e\int du\,{\cal J}\Phi^{(0)}\big|=$ $0\,$, (44d)
we can represent the variations of $S_{0}$, $S_{1}$ and $S_{2}$ in the forms
$\delta_{\xi}S_{0}=$ $\int d^{4}x\,e\int du(-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}$
$-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}})({\cal D}^{-})^{4}{\cal L}^{(+4)}\big|.$ (45a)
$\delta_{\xi}S_{1}=$ $\int d^{4}x\,e\int du\,S^{--}(-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}$
$+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}})[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}]{\cal L}^{(+4)}\big|\,$, (45b)
$\delta_{\xi}S_{2}=$ $\int d^{4}x\,e\int du(S^{--})^{2}(\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}-\xi^{+\alpha}{\cal D}^{-}_{\alpha}){\cal L}^{(+4)}\big|\,.$ (45c)
We will proceed by the following steps:
@@
    Commute the derivatives ${\cal D}^{+}_{\alpha}$ and $\bar{\cal D}^{+\dot{\alpha}}$ with $({\cal D}^{-})^{2}$ and $(\bar{\cal D}^{-})^{2}$ using the algebra of covariant derivatives (28) and the analyticity of the Lagrangian, ${\cal D}^{+}_{\alpha}{\cal L}^{(+4)}=\bar{\cal D}^{+\dot{\alpha}}{\cal L}^{(+4)}=0$.
@@
    Using the same algebra of covariant derivatives, in the integrand pull the operators ${\cal D}_{a}$ and $\cal J$ to the left and use the rules of integration by parts (44).
@@
    $\sigma^{a}_{\alpha\dot{\alpha}}{\cal D}_{a}\xi^{-\alpha}=$ $2i(S^{--}\bar{\xi}^{+}_{\dot{\alpha}}-S^{+-}\bar{\xi}^{-}_{\dot{\alpha}})\,$, (46)
    $(\sigma^{a})^{\alpha\dot{\alpha}}{\cal D}_{a}\bar{\xi}^{-}_{\dot{\alpha}}=$ $-2i(S^{--}\xi^{+\alpha}-S^{+-}\xi^{-\alpha})\,.$ (47)
    These identities follow from Eq. (11c).
As a result, we find the following expressions for the variations (45):
$\delta_{\xi}S_{0}=$ $\frac{1}{2}\int d^{4}x\,e\int du\,S^{--}\Big[\xi^{+\alpha}{\cal D}^{-}_{\alpha}(\bar{\cal D}^{-})^{2}$
$-\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}({\cal D}^{-})^{2}-8S^{--}\xi^{+\alpha}{\cal D}^{-}_{\alpha}+8S^{--}\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}$
$+12S^{+-}\xi^{-\alpha}{\cal D}^{-}_{\alpha}-12S^{+-}\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\Big]{\cal L}^{(+4)}\big|\,$, (48)
$\delta_{\xi}S_{1}=$ $-\int d^{4}x\,e\int du\,S^{--}\Big[\xi^{+\alpha}{\cal D}^{-}_{\alpha}(\bar{\cal D}^{-})^{2}$
$-\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}({\cal D}^{-})^{2}-8S^{--}\xi^{+\alpha}{\cal D}^{-}_{\alpha}+8S^{--}\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}$
$+12S^{+-}\xi^{-\alpha}{\cal D}^{-}_{\alpha}-12S^{+-}\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}\Big]{\cal L}^{(+4)}\big|\,.$ (49)
Hence, for $a_{1}=1/2$ and $a_{2}=0$, the action (41) is invariant under the isometry of AdS4|8, $\delta_{\xi}(S_{0}+\frac{1}{2}S_{1})=0$. Finally, we write the action principle for an analytic Lagrangian ${\cal L}^{(+4)}$ in the form:
$S=$ $\int d^{4}x\,e\int du\bigg\{({\cal D}^{-})^{4}$
$+\frac{1}{2}S^{--}\left[({\cal D}^{-})^{2}+(\bar{\cal D}^{-})^{2}\right]\bigg\}{\cal L}^{(+4)}\Big|\,.$ (50)
In this form, the analytic action principle is free from harmonic singularities and is suitable for studying the component structure of the analytic Lagrangian ${\cal L}^{(+4)}$. The relation of this action to Eq. (36) will be studied elsewhere.
Below, we demonstrate how the analytic action principle (50) applies for calculation of the component structure in the free $q$-hypermultiplet model.
V Example: component structure of the free hypermultiplet action
@@
${\cal L}^{(+4)}=-\tilde{q}^{+}D^{++}q^{+}\,$, (51)
where $q^{+}$ is an analytic superfield and $\tilde{q}^{+}$ is its tilde-conjugate,
${\cal D}^{+}_{\alpha}q^{+}=\bar{\cal D}^{+}_{\dot{\alpha}}q^{+}=0\,.$ (52)
Note that this superfield is not required to be annihilated by $D^{++}$ and, thus, is, in general, represented by an infinite series over the harmonic variables. This feature is crucial for the off-shell description of the $q$-hypermultiplet model in the harmonic superspace.
To evaluate the component structure of the action (50) with the Lagrangian (51), we proceed by the following steps:
@@
    We commute the derivatives $\mathcal{D}^{-}_{\alpha}$ and $\bar{\mathcal{D}}^{-\dot{\alpha}}$ to the right and act on $q^{+}$ and $\tilde{q}^{+}$. The commutators ${\cal D}^{-}_{\alpha},D^{++}]=-{\cal D}^{+}_{\alpha}$ and $\bar{\cal D}^{-\dot{\alpha}},D^{++}]=-\bar{\cal D}^{+\dot{\alpha}}$ yields the derivatives $\mathcal{D}^{+}$’s, which annihilate the analytic superfield $q^{+}$ according to Eq. (52).
@@
    The operators ${\cal D}_{a}$ and $\cal J$ are moved to the left by using the rules for integration by parts (44).
@@
$S_{q}=$ $S_{\mathrm{bos}}+S_{\mathrm{ferm}}\,$, (53)
$S_{\mathrm{bos}}=$ $-\int d^{4}x\,e\int du\Bigl[\bigl((\mathcal{D}^{-})^{4}\tilde{q}^{+}\bigr)D^{++}q^{+}+\tilde{q}^{+}D^{++}\bigl((\mathcal{D}^{-})^{4}q^{+}\bigr)+\frac{1}{4}(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}\tilde{q}^{+})D^{++}(\mathcal{D}^{-}_{\beta}\bar{\mathcal{D}}^{-}_{\dot{\alpha}}q^{+})$
@@
$+\frac{i}{2}(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}\tilde{q}^{+})(\mathcal{D}_{\beta\dot{\alpha}}q^{+})-\frac{i}{2}(\mathcal{D}_{\beta\dot{\alpha}}\tilde{q}^{+})(\mathcal{D}^{-\beta}\bar{\mathcal{D}}^{-\dot{\alpha}}q^{+})+\frac{1}{4}\bigl((\mathcal{D}^{-})^{2}\tilde{q}^{+}+(\bar{\mathcal{D}}^{-})^{2}\tilde{q}^{+}\bigr)(\mathcal{J}+2S^{--}D^{++})q^{+}$
$+\frac{1}{4}\tilde{q}^{+}(\mathcal{J}+2D^{++}S^{--})\bigl((\mathcal{D}^{-})^{2}q^{+}+(\bar{\mathcal{D}}^{-})^{2}q^{+}\bigr)+9S^{--}\tilde{q}^{+}\mathcal{J}q^{+}\Bigr]\bigg|\,$, (54)
$S_{\mathrm{ferm}}=$ $-\int d^{4}x\,e\int du\Bigl[\frac{1}{8}(\mathcal{D}^{-\alpha}(\bar{\mathcal{D}}^{-})^{2}\tilde{q}^{+})D^{++}(\mathcal{D}^{-}_{\alpha}q^{+})+\frac{1}{8}(\bar{\mathcal{D}}^{-}_{\dot{\beta}}\tilde{q}^{+})D^{++}(\bar{\mathcal{D}}^{-\dot{\beta}}(\mathcal{D}^{-})^{2}q^{+})$
@@
$+\frac{i}{2}(\mathcal{D}^{-\alpha}\tilde{q}^{+})(\mathcal{D}_{\alpha\dot{\beta}}\bar{\mathcal{D}}^{-\dot{\beta}}q^{+})-\frac{i}{2}(\bar{\mathcal{D}}^{-\dot{\beta}}\tilde{q}^{+})(\mathcal{D}_{\alpha\dot{\beta}}\mathcal{D}^{-\alpha}q^{+})+\frac{1}{2}(\mathcal{D}^{-\alpha}\tilde{q}^{+})(\mathcal{J}+2S^{+-}+2S^{--}D^{++})(\mathcal{D}^{-}_{\alpha}q^{+})$
$+\frac{1}{2}(\bar{\mathcal{D}}^{-}_{\dot{\beta}}\tilde{q}^{+})(\mathcal{J}+2S^{+-}+2S^{--}D^{++})(\bar{\mathcal{D}}^{-\dot{\beta}}q^{+})\Bigr]\bigg|\,.$ (55)
@@
These components enter the actions (54) and (55) as follows:
@@
$-\frac{1}{2}\Phi^{\alpha}({\cal J}+2S^{+-}+2S^{--}D^{++})\Psi_{\alpha}+\frac{1}{2}\bar{\Psi}_{\dot{\alpha}}({\cal J}+2S^{+-}+2S^{--}D^{++})\bar{\Phi}^{\dot{\alpha}}\Big].$ (57b)